OpenStudy (anonymous):

Match each description when z = 9 + 3i. 1. Real part of z a.3i 2. Imaginary part of b.z9 3. Complex conjugate of c.z9 - 3i 4. 3i - z d.-3i 5. z - 9 e.3 6. 9 - z f.-9

OpenStudy (anonymous):

just treat them separately like apples and oranges 9=apples 3i oranges z=9+3i ==real part +imaginary part 1)9 2)3i 3)(9-3i) change of sign of given(9+3i_) 4) 3i-z =3i -9-3i =-9 5) z-9 =9+3i-9 =3i 6) 9-z =9-9-3i =- 3i just treat them seperately like apples and oranges 9 =apples 3i =oranges